Translate the following situation into an equation. Do not solve.
“Profits in the fourth quarter for Company A were 78.2% of that of Company B. Profits in the fourth quarter for Company C were 103.8% of Company B’s profits. Together the three companies earned $324,300 in the fourth quarter. What were the profits for Company B?”

Answer by robertb(5830)

A = 0782*B,
C = 1.038*B,
A+B+C = 324,300.